the number of homes built can be modeled by an exponential function,
H= p * at , where p is the number of new homes built in the first year recorded.
p=10 t =3 a=2
I'm not sure how to calculate

Answer by rfer(10417)

H=p*at is the same as H=pat
H=10(2)(3)
H=60